OpenStudy (anonymous):

find g(f(x)) if f(x)=3x-2 g(x)= x+2

OpenStudy (anonymous):

x-2(3x-2(x)) =x-2(3x^2-2x) =3x^3-2x^2+6x^2-4x =3x^3+4x^2-4x

OpenStudy (whpalmer4):

basically, you take the equation for f(x), and substitute the right hand side wherever you see x in g(x). f(x) = 3x-2 g(x) = x+2 g(f(x)) = (3x-2) + 2

OpenStudy (anonymous):

f(x)=3x-2 g(x)=x+2 g(f(x))=(3x-2)+2 =3x-2+2 =3x